# KAT-Coder Technical Report

### v1

@@ Abstract
Recent advances in large language models (LLMs) have enabled progress in agentic coding, where models autonomously reason, plan, and act within interactive software development workflows[1, 2, 3, 4, 5, 6, 7, 8]. However, bridging the gap between static text-based training and dynamic real-world agentic execution remains a core challenge. In this technical report, we present KAT-Coder, a large-scale agentic code model trained through a multi-stage curriculum encompassing Mid-Term Training, Supervised Fine-Tuning (SFT), Reinforcement Fine-Tuning (RFT), and Reinforcement-to-Deployment Adaptation. The Mid-Term stage enhances reasoning, planning, and reflection capabilities through a corpus of real software engineering data and synthetic agentic interactions. The SFT stage constructs a million-sample dataset balancing twenty programming languages, ten development contexts, and ten task archetypes. The RFT stage introduces a novel multi-ground-truth reward formulation for stable and sample-efficient policy optimization. Finally, the Reinforcement-to-Deployment phase adapts the model to production-grade IDE environments using Error-Masked SFT and Tree-Structured Trajectory Training[9]. In summary, these stages enable KAT-Coder to achieve robust tool-use reliability, instruction alignment, and long-context reasoning, forming a deployable foundation for real-world intelligent coding agents. Our KAT series 32B model, KAT-Dev, has been open-sourced on [Uncaptioned image] https://huggingface.co/Kwaipilot/KAT-Dev.
@@ 1 Introduction
Recent advances in Large Language Models (LLMs) have catalyzed a shift from static text generation toward agentic intelligence, where models autonomously reason, plan, and act within dynamic environments. In software engineering, this transformation manifests as agentic coding—a paradigm in which models function as collaborative problem solvers rather than passive code generators. Despite rapid progress, a fundamental challenge persists: bridging the gap between static, text-based training and interactive, real-world execution. Conventional code models, typically trained on massive but inert text corpora, lack the adaptive reasoning and contextual control required to operate reliably in live integrated development environments (IDEs).

### v2

@@ Abstract
Recent advances in large language models (LLMs) have enabled progress in agentic coding, where models autonomously reason, plan, and act within interactive software development workflows. However, bridging the gap between static text-based training and dynamic real-world agentic execution remains a core challenge. In this technical report, we present KAT-Coder, a large-scale agentic code model trained through a multi-stage curriculum encompassing Mid-Term Training, Supervised Fine-Tuning (SFT), Reinforcement Fine-Tuning (RFT), and Reinforcement-to-Deployment Adaptation. The Mid-Term stage enhances reasoning, planning, and reflection capabilities through a corpus of real software engineering data and synthetic agentic interactions. The SFT stage constructs a million-sample dataset balancing twenty programming languages, ten development contexts, and ten task archetypes. The RFT stage introduces a novel multi-ground-truth reward formulation for stable and sample-efficient policy optimization. Finally, the Reinforcement-to-Deployment phase adapts the model to production-grade IDE environments using Error-Masked SFT and Tree-Structured Trajectory Training. In summary, these stages enable KAT-Coder to achieve robust tool-use reliability, instruction alignment, and long-context reasoning, forming a deployable foundation for real-world intelligent coding agents. Our KAT series 32B model, KAT-Dev, has been open-sourced on [Uncaptioned image] https://huggingface.co/Kwaipilot/KAT-Dev.
@@ 1 Introduction
Recent advances in Large Language Models (LLMs) have catalyzed a shift from static text generation toward agentic intelligence, where models autonomously reason, plan, and act within dynamic environments. In software engineering, this transformation manifests as agentic coding—a paradigm in which models function as collaborative problem solvers rather than passive code generators. Despite rapid progress, a fundamental challenge persists: bridging the gap between static, text-based training and interactive, real-world execution. Conventional code models, typically trained on massive but inert text corpora, lack the adaptive reasoning and contextual control required to operate reliably in live integrated development environments (IDEs) [1, 2, 3, 4, 5, 6, 7, 8].